### v5

# UFO [Uncaptioned image]: A UI-Focused Agent for Windows OS Interaction

Abstract
We introduce UFO [Uncaptioned image], an innovative UI-Focused agent to fulfill user requests tailored to applications on Windows OS, harnessing the capabilities of GPT-Vision. UFO employs a dual-agent framework to meticulously observe and analyze the graphical user interface (GUI) and control information of Windows applications. This enables the agent to seamlessly navigate and operate within individual applications and across them to fulfill user requests, even when spanning multiple applications. The framework incorporates a control interaction module, facilitating action grounding without human intervention and enabling fully automated execution. Consequently, UFO transforms arduous and time-consuming processes into simple tasks achievable solely through natural language commands. We conducted testing of UFO across 9 popular Windows applications, encompassing a variety of scenarios reflective of users’ daily usage. The results, derived from both quantitative metrics and real-case studies, underscore the superior effectiveness of UFO in fulfilling user requests. To the best of our knowledge, UFO stands as the first UI agent specifically tailored for task completion within the Windows OS environment. The open-source code for UFO is available on https://github.com/microsoft/UFO.
Figure 1: An illustration of the high-level concept of the Windows agent UFO. It completes a user request by extracting information from a Word document, observing a figure in Photos, and summarizing content from a PowerPoint presentation on Windows OS. Then, UFO utilizes this information to compose an email and send, all accomplished automatically.
1 Introduction
The advent of Large Language Models (LLMs) Ouyang et al. (2022); Achiam et al. (2023); Touvron et al. (2023) has showcased revolutionary potential in solving complex problems akin to human reasoning, planning, and collaboration Hao et al. (2023); Ding et al. (2023); Talebirad & Nadiri (2023). This development brings humanity closer to the realization of Artificial General Intelligence (AGI) Liu et al. (2023), offering assistance in various tasks in our daily lives and exhibiting a robust and comprehensive level of ability that was previously considered challenging Chen et al. (2023b); Shao et al. (2023); Jin et al. (2023). The advancement towards more sophisticated multi-modal visual LLMs, exemplified by GPT-Vision Zhang et al. (2023); Yang et al. (2023c), introduces a visual dimension to LLMs, expanding their capabilities to encompass a myriad of visual tasks and broadening their scope to address challenges that require visual capability Wu et al. (2023a); Zheng et al. (2024); Wake et al. (2023).
The applications of Visual Large Language Models (VLM) are continually emerging and thriving. One notable application involves using VLMs to interact with the User Interface (UI) or Graphical User Interface (GUI) of software applications, fulfilling user requests expressed in natural language and grounding them in physical devices. While GUIs are primarily designed for human users to see and interact with, the elements and controls within the application’s UI provide a crucial bridge for VLMs to interact, especially by perceiving their visual information Hong et al. (2023b) and grounding their actions in a manner similar to humans. This facilitates the evolution from Large Language Models (LLMs) to Large Action Models (LAMs) Brohan et al. (2023), enabling their decisions to manifest in physical actions and have tangible real-world impacts.
In this context, the Windows operating system (OS) stands out as representative platform for LAMs, due to its high market share in the daily use of computer systems Adekotujo et al. (2020), the presence of versatile applications and GUIs built upon it Ramler et al. (2018), and the complexity of tasks that necessitate long-term planning and interaction across various applications Stallings (2005). The prospect of having a general intelligent agent that can comprehend user requests in natural language, and autonomously interact with the UIs of applications built on Windows is highly appealing. Despite the evident demand for developing VLM agents tailored for Windows OS to fulfill user requests, this direction remains largely unexplored, as most existing agents predominantly focus on smartphones Yang et al. (2023b) or web applications Zheng et al. (2024). This gap in exploration presents a vast, untapped opportunity to develop a Windows-specific agent.
To bridge this gap, we introduce UFO [Uncaptioned image], a specialized UI-Focused Agent designed for seamless interaction with the Windows OS, leveraging the cutting-edge capabilities of the VLM GPT-Vision Yang et al. (2023c). UFO adopts a dual-agent framework, with each agent analyzing screenshots and extracting information from the GUI to make informed decisions in selecting applications. Subsequently, it navigates and executes actions on their controls, emulating human users to fulfill requests in natural language. The system incorporates a control interaction component, which plays a crucial role in translating actions from GPT-Vision into grounded execution on applications. This feature ensures complete automation without the need for human intervention, thereby establishing it as a comprehensive LAM framework.
Acknowledging that user requests often span multiple applications in their daily routines, UFO incorporates an application-switching mechanism, allowing it to transition to a different application when needed. This expanded capability equips UFO to handle more intricate tasks that are typically overlooked by other agents. Furthermore, UFO is highly extensible, enabling users to design and customize actions and controls for specific tasks and applications, enhancing its versatility. In summary, UFO streamlines various tasks for users engaged in daily computer activities, transforming lengthy and tedious processes into simple tasks achievable through only textual commands. This positions UFO as a valuable, user-friendly, and automated co-pilot for the Windows OS, effectively reducing the overall complexity of usage. We illustrate this concept at a high level in Figure 1, where UFO composes and sends an email by integrating text extracted from a Word document, observations from an image, and summaries from a PowerPoint presentation.
In order to assess its efficacy, we conducted versatile testing of our UFO framework, subjecting it to a battery of 50 tasks spanning 9 widely used Windows applications. These tasks were carefully chosen to cover a diverse spectrum of scenarios reflective of users’ daily computational needs. The evaluation involved both quantitative metrics and in-depth case studies, highlighting the robustness and adaptability of our design, especially in the context of extended and intricate requests that traverse multiple applications. To the best of our knowledge, UFO stands as the pioneering agent tailored for general applications within the Windows OS environment.
2 Related Work
In this section, we review research efforts relevant to UFO, with a focus on the domains of LLM agents and LLM-based GUI intelligence.
2.1 LLM Agents
The advent of LLM agents has significantly expanded the capabilities of LLMs Wang et al. (2023); Xi et al. (2023); Talebirad & Nadiri (2023), empowering them to engage in planning, observation, memorization, and responsive actions. This augmentation enables LLMs to undertake more intricate tasks by emulating human-like decision-making processes. Notably, AutoGPT Significant Gravitas stands as a pioneering agent within this domain, facilitating interaction with users and decomposing LLMs’ actions into discrete components such as thoughts, reasoning, and criticism, thereby effectively addressing user requests. Additionally, TaskWeaver Qiao et al. (2023) represents a noteworthy code-centric agent framework designed to disassemble user requests and translate them into manageable subtasks executable via Python code. The LangChain Agent Chase (2022), serves as an extension of the LangChain framework. This extension empowers the utilization of a LLM to intelligently select a sequence of actions, which may involve the utilization of customized tools. The integration of LLMs within such agents not only augments their decision-making capabilities but also signifies a significant stride towards the eventual realization of AGI.
Furthermore, the incorporation of multi-agent LLMs represents a more potent and scalable framework. This architecture facilitates the allocation of tasks to individual agents based on their respective strengths, fostering collaboration or competition among agents to effectively accomplish complex tasks. AutoGen Wu et al. (2023b) exemplifies this approach by designing each agent to be highly customizable and conversable. This design philosophy enables each agent to leverage its specific strengths, thereby contributing optimally to the overall task accomplishment within the multi-agent system. MetaGPT Hong et al. (2023a) constitutes another notable multi-agent framework. Within this framework, distinct roles are assigned to individual GPTs, collectively forming a collaborative software entity adept at addressing complex tasks. Furthermore, another multi-agent framework named AutoAgents Chen et al. (2023a) generates and coordinates multiple specialized agents to form an AI team tailored for complex tasks. These multi-agent frameworks signify a burgeoning branch in LLM agent development, offering heightened scalability for tackling intricate tasks.
2.2 LLM-based GUI Intelligence
The utilization of multimodal LLM systems Durante et al. (2024) for navigating and controlling GUIs in applications has emerged as a prominent and burgeoning research area. Yan et al., Yan et al. (2023) employed GPT-4V Yang et al. (2023c) to navigate mobile applications by inputting screenshots of GUIs to the LLM, demonstrating state-of-the-art performance across various datasets and human-designed navigation tasks. Similarly, AppAgent yang2023HostAgent leverages GPT-4V as smartphone users, enabling them to take actions on mobile applications based on smartphone snapshots, thereby autonomously fulfilling user requests on physical phones. MobileAgent Wang et al. (2024) integrates Optical Character Recognition (OCR) tools to augment the GPT-V employed within a comparable mobile agent designed for task completion on mobile phones. This integration results in a notable enhancement, enabling MobileAgent to achieve completion rates comparable to human performance. On the other hand, CogAgent Hong et al. (2023b) takes a distinct approach by training a dedicated visual language model specialized in GUI understanding and navigation, providing a more tailored solution for various GUI intelligence tasks.
Distinguishing itself from existing frameworks, our proposed UFO stands out as a multimodal LLM agent framework specifically tailored to fulfilling user requests and manipulating application UI within the Windows OS. This framework transcends the constrains posed by different applications, enabling the seamless completion of more intricate tasks within the Windows environment.
3 The Design of UFO
We present UFO [Uncaptioned image], a groundbreaking UI-Focused Multimodal Agent designed for the Windows OS interaction. UFO possesses the capability to comprehend users’ requests expressed in natural language, breaking them down into a series of step-by-step sub-tasks. It then observe the UI screenshots of of applications and operates on their control elements to fulfill the overall objective. This unique functionality enables UFO to seamlessly navigate across multiple applications, effectively completing complex tasks and transcending the boundaries of different applications. In Section 3.1, we provide an overview of the design of UFO, followed by detailed discussions on each core component in the subsequent subsections.
3.1 UFO in a Nutshell
Figure 2: The overall architecture of the UFO.
First, we present the comprehensive architecture of UFO in Figure 2. UFO operates as a dual-agent framework, encompassing (i) an HostAgent tasked with choosing an application for fulfilling user requests. This agent may also switch to a different application when a request spans multiple applications, and the task is partially completed in the preceding application. Additionally, (ii) an AppAgent is incorporated, responsible for iteratively executing actions on the selected applications until the task is successfully concluded within a specific application. Both agents leverage the multi-modal capabilities of GPT-Vision to comprehend the application UI and fulfill the user’s request. They utilize a Control Interaction module to ground their actions, enabling tangible impacts on the system.
Upon receiving a user request, the HostAgent undertakes an analysis of the demand. It endeavors to select a fitting application for fulfilling the request from the currently active applications. UFO equips HostAgent with the full desktop screenshot and a list of available applications for reference, facilitating HostAgent’s decision-making process. Subsequently, HostAgent selects an appropriate application and formulates a comprehensive global plan for request completion. This plan is then passed on to the AppAgent.
Once a suitable application is identified, it is brought into focus on the desktop. The AppAgent then initiates actions to fulfill the user request. Before each action selection step, UFO captures screenshots of the current application’s UI window with all available controls annotated. Additionally, UFO records information about each control for AppAgent’s observation. AppAgent is tasked with choosing a control to operate and subsequently selecting a specific action to execute on the chosen control via a control interaction module. This decision is based on AppAgent’s observation, its prior plan, and its operation memory. Following the execution, UFO constructs a local plan for future steps, and proceeds to the next action selection step. This recursive process continues until the user request is successfully completed within the selected application. This concludes one phase of the user request.
In scenarios where the user request spans multiple applications, the AppAgent will delegate the task to the HostAgent for the purpose of switching to a different application once AppAgent completes its tasks on the current one, initiating the second phase of the request. This iterative process continues until all aspects of the user request are fully completed. Users have the option to introduce new requests interactively, prompting UFO to process the new request by repeating the aforementioned process. Upon the successful completion of all user requests, UFO concludes its operation. In the subsequent sections, we delve into the intricate details of each component within the UFO framework.
3.2 HostAgent
Figure 3: An illustration of the HostAgent.
The HostAgent bears the responsibility of selecting an active application to fulfill user requests or switching to a new application when necessary. Additionally, HostAgent constructs a comprehensive global plan to orchestrate the entire task, and its architecture is illustrated in Figure 3. HostAgent takes the following information as input:
  • User Request: The original user query submitted to UFO.
  • Desktop Screenshots: Screenshots of the current desktop, where multiple screens are concatenated into a single image.
  • Application Information: A listing of available active application details, encompassing their names and types.
  • Memory: Comprising previous thoughts, comments, actions, and execution results.
  • Examples: Textual examples for application selection, serving as demonstrations for the task.
The provided information, including Desktop Screenshots, Application Information, and Memory, collectively equips HostAgent with a comprehensive set of data to facilitate decision-making. Desktop Screenshots and Application Information enable HostAgent to comprehend the current status and constrain its scope for application selection. On the other hand, Memory serves as a historical record of past request completions, aiding HostAgent in making informed decisions based on prior experiences. This multifaceted input framework enhances HostAgent’s ability to select the most suitable application for fulfilling user requests.
Upon collecting all pertinent information, HostAgent employs GPT-V to generate the following outputs:
  • Observation: Detailed descriptions of the screenshots of the current desktop window.
  • Thoughts: The logical next step required to fulfill the given task, adhering to the Chain-of-thought (CoT) paradigm Wei et al. (2022).
  • Selected Application: The label and name of the chosen application.
  • Status: The task status, denoted as either “CONTINUE” or “FINISH”.
  • Global Plan: A subsequent plan of action to complete the user request, typically a global and coarse-grained plan.
  • Comment: Additional comments or information to provide, including a brief progress summary and points to be highlighted.
Prompting HostAgent to provide its Observation and Thoughts serves a dual purpose. Firstly, it encourages HostAgent to meticulously analyze the current status, offering a detailed explanation for its logic and decision-making process. This not only enhances the logical coherence of its decisions Wei et al. (2022); Ding et al. (2023), but also contributes to improving the overall interpretability of UFO. Secondly, HostAgent determines the status of the task, outputting “FINISH” if it deems the task completed. HostAgent may also leave comments for the user, reporting progress, highlighting potential issues, or addressing any user queries. Once HostAgent identifies the selected application, UFO proceeds to take specific actions within this application to fulfill the user request, with the AppAgent responsible for executing these actions.
3.3 AppAgent
Figure 4: An illustration of the AppAgent.
The AppAgent functions as a downstream entity following HostAgent, responsible for executing specific actions on the selected application to fulfill the user request. Its input and output also exhibit variations compared to HostAgent, as depicted in Figure 4. AppAgent accepts the following input:
  • User Request: The original user query submitted to UFO, identical to the HostAgent.
  • Screenshots: The screenshots encompass three parts: (i) Previous Screenshot; (ii) Clean screenshot; and (iii) Annotated screenshot.
  • Control Information: A listing of names and types of controls that are enabled for operations in the selected application.
  • Memory: Previous thoughts, comments, actions, and execution results, mirroring the HostAgent.
  • Examples: Textual examples for action selection, serving as demonstrations.
In contrast to the HostAgent, UFO provides AppAgent with three types of screenshots to aid in its decision-making process. The previous screenshot with the last selected control highlighted in a red rectangle (i.e.,  ) helps comprehend the operation execution in the last step and analyze the impact of the action. The clean screenshot allows for an understanding of the application’s status without annotation obstructions, and the annotated screenshot, labeling each control with a number (e.g., 36) using Set-of-Mark (SoM) Yang et al. (2023a), facilitates a better understanding of the function and location of UI elements. Different types of controls are labeled with distinct colors for differentiation.
Moreover, the memory fed into AppAgent serves two key purposes. Firstly, it acts as a reminder for the agent, enabling AppAgent to analyze past actions and reduce the likelihood of repeating actions that proved ineffective. Secondly, it establishes a crucial channel for cross-application communication. Execution results, such as text extracted from a document or a description of an image, are stored in the memory module. AppAgent can selectively incorporate this information for actions that require it, such as composing an email with text from different sources. This augmentation significantly extends the capabilities of UFO.
Given this comprehensive input, AppAgent meticulously analyzes all the information and outputs the following:
  • Observation: Detailed descriptions of the screenshots of the current application window, along with an analysis of whether the last action has taken effect.
  • Thoughts: The logical thinking and rationale process behind the current action decision.
  • Selected Control: The label and name of the chosen control for the operation.
  • Function: The specific function and its arguments applied to the control.
  • Status: The task status, indicated as either “CONTINUE” if further action is needed, “FINISH” if the task is completed, “PENDING” if the current action requires user confirmation, “SCREENSHOT” if the agent believes a further screenshot is needed to annotate a smaller set of controls, and “APP_SELECTION” when the task is completed on the current application and a switch to a different one is required.
  • Local Plan: A more precise and fine-grained plan for future actions to completely fulfill the user request.
  • Comment: Additional comments or information, including a brief progress summary, highlighted points, or changes in the plan, similar to what HostAgent provides.
While some of its output fields may share similarities with HostAgent, UFO determines the next step based on the outputted Status of the task. If the task is not finished, it applies the Function to the selected control, triggering the next state of the application after execution. AppAgent iteratively repeats this process of observing and reacting to the selected application until the user request is fully completed or a switch to a different application is required.
3.4 Control Interaction
Figure 5: An example of the annotated PowerPoint GUI with information provided by pywinauto. Different colors of annotations represent different control types.
To execute and ground the selected action from AppAgent on the application controls, UFO first detects and translates the action into an executable operation. The Python package pywinauto Bim & Min-shuai (2014) offers valuable tools for inspecting UI controls and performing actions on them. It provides a range of functions and classes for automating the Windows GUI, allowing for programmatic interaction and control of Windows applications. This capability is particularly useful for tasks such as automated testing, scripting, and the automation of repetitive tasks. As the backend, we choose the Windows UI Automation API Dinh et al. (2018) for its robust support in UI inspection and interaction through code.
UFO utilizes pywinauto to inspect all actionable controls of an application, retrieving their precise location and boundary box to facilitate annotations with SoM Yang et al. (2023a). An example of annotated controls on a PowerPoint GUI is presented in Figure 5, with control information provided by pywinauto. Different colors various control types. Additionally, pywinauto provides rich context for each control, including its name, type, and title, which are crucial information for control and action selection.
Table 1: The detailed descriptions of control types supported by UFO.
Control Type Description
Button A button is a user interface element that users can interact with to trigger an action. Clicking the button typically initiates a specific operation or command.
Edit An edit control allows users to input and edit text or numeric data. It is commonly used for fields where users can type information, such as textboxes or search bars.
TabItem A tab item is part of a tab control, organizing content into multiple pages. Users can switch between different tab items to access distinct sets of information or functionalities.
Document A document control represents a document or a page in a document-view architecture. It is often used to display and manage documents or large blocks of text.
ListItem A list item is an element within a list control, presenting data in a list format. Users can select and interact with individual items within the list.
MenuItem A menu item is part of a menu control, providing a list of commands or options. Users can click on menu items to trigger specific actions or navigate through application features.
TreeItem A tree item is a node within a tree control, organizing information in a hierarchical structure. Users can expand or collapse tree items to navigate through a hierarchical set of data.
ComboBox A ComboBox is a combination of a text box and a drop-down list. It allows users to either type a value directly into the text box or choose from a predefined list by opening the drop-down menu.
Hyperlink A Hyperlink enables users to navigate to another location or resource. They are often used to provide easy access to external websites, documents, or specific sections within an application.
ScrollBar A scroll bar allows users to scroll through content that is larger than the visible area. It provides a way to navigate vertically or horizontally within a window or control.
UFO focuses on the following 10 constrained control types with high relevance, as determined by our analysis. These include Button, Edit, TabItem, Document, ListItem, MenuItem, TreeItem, ComboBox, Hyperlink, ScrollBar. We show a detailed descriptions of these control types in Table 1. These set can covers majority of relevant controls in the applications, and it is also extendable per request.
For the specific functions applied to the control, we have chosen common and widely used mouse operations supported by pywinauto, as well as developed customized actions. These actions include [^1] Note that each of these actions can only be executed on specific and different types of controls.:
  • Click: Clicking the control item with the mouse, with options for left or right clicks, single or double clicks.
  • SetText: Inputting text into an editable control, mimicking the keyboard behaviors.
  • Annotate: Capturing a screenshot of the current application window and annotating the control item on the GUI.
  • Summary: Summarizing the observation of the current application window based on the clean screenshot.
  • GetText: Retrieving the textual information of the control.
  • Scroll: Scrolling the control item vertically or horizontally to make hidden content visible.
Click, SetText, GetText, and Scroll are common functions originally supported by pywinauto, covering most daily operations on GUI. Annotate and Summary are customized operations to fulfill special requests for UFO. The former allows re-annotation of the GUI with a more concise list of controls (details in Section 3.5.3), and the latter enables UFO to describe its visual observation in text to fulfill user requests. At every step, AppAgent will choose an action from this available list to execute on the selected UI control in the application. With the control interaction module, UFO evolves into a LAM capable of making tangible impacts on the system.
3.5 Special Design Consideration
UFO incorporates a series of dedicated design elements specifically tailored to the Windows OS. These enhancements are aimed at facilitating more effective, automated, and secure interactions with UI controls, thereby enhancing its capacity to address users’ requests. Key features encompass interactive mode, action customization, control filtering, plan reflection, and safeguard, each of which is elaborated upon in the subsequent subsections.
3.5.1 Interactive Mode
UFO offers users the capability to engage in interactive and iterative interactions, rather than insisting on one-shot completions. Upon the completion of a task, users have the flexibility to request UFO to enhance the previous task, propose an entirely new task for UFO to undertake, or perform operations to assist UFO in tasks where it may lack proficiency, such as providing a password input. This user-friendly approach not only distinguishes UFO from other existing UI agents in the market but also allows it to absorb user feedback, facilitating the completion of longer and more intricate tasks.
3.5.2 Action Customization
UFO currently facilitates operations on controls or the UI, as detailed in Section 3.4. However, it is imperative to note that this list is not exhaustive and can be highly extended and customized beyond the constraints of Windows UI Automation. This extensibility is crucial for tailoring the framework to meet specific user needs, allowing for the incorporation of functionalities such as keyboard shortcuts, macro commands, plugins, and more. An illustrative example is the utilization of summary(), which leverages UFO’s multimodal capabilities to observe and describe screenshots as required.
To achieve this level of customization, users can register their bespoke operations. This involves specifying the purpose, arguments, return values, and, if necessary, providing illustrative examples for demonstration purposes. This information is then incorporated into the prompt for UFO’s reference. Once the registration process is completed, the customized operation becomes available for execution by UFO. This inherent flexibility renders UFO a highly extendable framework, enabling it to fulfill more intricate and user-specific requests within the Windows system.
3.5.3 Control Filtering
In the GUI of an application, Windows UI Automation can detect hundreds of control items, each available for operations. However, annotating all these controls can clutter the application UI screenshots, obstructing the view of individual items and generating an extensive list that may pose challenges for UFO in making optimal choice decisions. Notably, some of these controls may prove unlikely to be useful or relevant for fulfilling user requests. Consequently, the implementation of filtering mechanisms becomes crucial to exclude certain controls and streamline the decision-making process for UFO.
To address this challenge, UFO employs a dual-level control filtering approach, comprising the hard level and the soft level. At the hard level, candidate controls are constrained based on specific control types with high relevance and popularity, as detailed in Section 3.4. Moreover, we incorporate a soft filter mechanism that empowers UFO to dynamically determine whether to re-select a more concise list of specified controls. This adaptive filtering approach is triggered when UFO perceives an excessive number of controls, potentially cluttering the current screenshot and obstructing the visibility of the required control. In such scenarios, UFO intelligently returns a refined list of control candidates of interest. Subsequently, it captures a new screenshot annotated only with these controls, facilitating a more focused and effective filtering process. This feature enhances the automation capabilities of the framework, enabling UFO to make intelligent decisions for optimal performance.
3.5.4 Plan Reflection
While both the application selection agent and the action selection agent are responsible for initiating plans to fulfill user requests, the actual state of the application’s UI may not always align with the anticipated conditions. For instance, if UFO initially plans to click a “New Email” button in the next step, but this button is not visible in the current UI, UFO may need to first navigate to the ”Main Page” and subsequently locate the “New Email” button. Consequently, both the plan and action should adapt accordingly.
To address this dynamic nature of the UI, we prompt UFO to continuously revise its plan at every decision step, allowing it to deviate from the original course as needed. This adaptive approach enhances UFO’s responsiveness to the evolving application status based on its observations. The efficacy of such reflective mechanisms has been substantiated in various LLM frameworks and agent architectures Qiao et al. (2023); Ding et al. (2023); Shinn et al. (2023). Moreover, the integration of plan reflection significantly contributes to the enhanced performance of UFO in navigating and interacting with diverse application UIs.
3.5.5 Safeguard
Lastly, we acknowledge the sensitivity of certain actions within the system, such as irreversible changes resulting from operations like file deletion. In recognition of these potential risks, UFO incorporates a safeguard mechanism to seek user confirmation before executing such actions. The safeguard functionality is not limited to the following enumerated list in Table 2, as UFO intelligently assesses the sensitivity of each action. With the deployment of this safeguard, UFO establishes itself as a significantly safer and trustworthy agent, mitigating the risk of compromising the system or jeopardizing user files and privacy.
Table 2: An incomplete list of sensitive actions considered in UFO.
Sensitive Action Description
Send Action for a Message or Email Initiating the “Send” action, such as clicking the send button, is considered sensitive as the dispatched message or email becomes irretrievable.
Deleting or Modifying Files and Folders Operations involving the deletion or modification of files and folders, especially those situated in critical system directories or containing vital user data.
Closing a Window or Application Closing a window or application is flagged as sensitive due to the potential for data loss or system crashes.
Accessing Webcam or Microphone Accessing the webcam or microphone without explicit user consent is identified as sensitive to address privacy concerns.
Installing or Uninstalling Software Actions related to installing or uninstalling software applications are marked as sensitive due to their impact on system configuration and potential security risks.
Browser History or Password Retrieval Retrieving sensitive user data such as browser history or stored passwords is identified as a sensitive action, posing potential privacy leaks.
4 Experiment
In this section, we comprehensively evaluate UFO’s performance in completing user requests on the Windows OS. The evaluation is conducted through a combination of quantitative analysis and case studies, encompassing diverse tasks.
4.1 Benchmark & Baselines & Metrics
To comprehensively evaluate the performance across various Windows applications, we developed a benchmark called WindowsBench. This benchmark comprises 50 user requests, encompassing 9 popular Windows applications commonly used in daily tasks. The selected applications include Outlook, Photos, PowerPoint, Word, Adobe Acrobat, File Explorer, Visual Studio Code, WeChat, and Edge Browser. These applications cater to different purposes such as work, communication, coding, reading, and web browsing, ensuring the diversity and comprehensiveness of the evaluation. For each application, we designed 5 distinct requests, and an additional 5 requests involve interactions spanning multiple applications. This setup results in a total of 50 requests, with each application having at least one request linked to a follow-up request, providing a comprehensive evaluation of UFO’s interactive mode. We present the detailed list of requests utilized in WindowsBench in Table 5, 6, and 7 in Appdendix Section A. Requests involving follow-up interactions are organized numerically within each category.
Given the absence of existing Windows agents, we selected GPT-3.5 and GPT-4 as baseline models. Since these models lack the capability to directly interact with applications, we instructed them to provide step-by-step instructions to complete the user requests. A human then acted as their surrogate to carry out the operations. When visual abilities were required, we allowed the baselines to pause, as they couldn’t perform these tasks independently.
In terms of evaluation metrics, we assess UFO from three perspectives for each request: success, step, completion rate, and safeguard rate. The success metric determines if the agent successfully completes the request. The step refers to the number of actions the agent takes to fulfill a task, serving as an indicator of efficiency. The completion rate is the ratio of the number of correct steps to the total number of steps. Lastly, the safeguard rate measures how often UFO requests user confirmation when the request involves sensitive actions. Given the potential instability of GPT-V, which may generate different outputs each time, we conduct three tests for each request and select the one with the highest completion rate. This approach is consistent for the other baselines as well.
4.2 Performance Evaluation
Table 3: Performance comparison achieved by UFO on WindowsBench.
Framework Success Step Completion Rate Safeguard Rate
GPT-3.5 (Human Surrogate) 24% 7.86 31.6% 50%
GPT-4 (Human Surrogate) 42% 8.44 47.8% 57.1%
UFO 86% 5.48 89.6% 85.7%
Table 4: The detailed performance breakdown across applications achieved by UFO on WindowsBench.
Application Success Step Completion Rate Safeguard Rate
Outlook 100.0% 6.8 94.0% 100.0%
Photos 80.0% 4.0 96.7% 100.0%
PowerPoint 80.0% 5.6 88.8% 50.0%
Word 100.0% 5.4 92.7% -
Adobe Acrobat 60.0% 4.2 78.7% 100.0%
File Explorer 100.0% 4.8 88.7% 100.0%
Visual Studio Code 80.0% 4.0 84.0% -
WeChat 100.0% 5.0 98.0% 66.7%
Edge Browser 80.0% 5.2 92.0% 100.0%
Cross-Application 80.0% 9.8 83.0% 100.0%
Firstly, we present a comprehensive quantitative comparison of various frameworks using the WindowsBench dataset, as illustrated in Table 3. Remarkably, our UFO achieves an impressive 86% success rate across the benchmark, surpassing the best-performing baseline (GPT-4) by more than double. This result underscores the sophistication of UFO in successfully executing tasks on the Windows OS, positioning it as a highly effective agent. Moreover, UFO exhibits the highest completion rate, suggesting its capability to take significantly more accurate actions. We also observe that UFO completes tasks with the fewest steps, showcasing its efficiency as a framework, while GPT-3.5 and GPT-4 tend to provide more steps but are less effective for the tasks. From the safety perspective, UFO attains the highest safeguard rate at 85.7%, which proves that it can classify sensitive requests accurately, affirming its status as a secure agent that actively seeks user confirmation for these requests.
The inferior performance of the baselines compared to UFO can be attributed to two primary factors. Firstly, both baselines lack the ability to directly interact with the real application environment, relying on human surrogates for action execution. This limitation results in an inability to adapt to changes and reflections in the environment, leading to decreased accuracy. Secondly, the baselines solely accept textual input, neglecting the importance of visual capabilities for GUI interaction. This weakness hinders their effectiveness in completing user requests on Windows, where visual information is often crucial. Notably, GPT-4 outperforms GPT-3.5, highlighting its greater potential in these tasks. In summary, considering all the results above, we show that UFO excels across all four evaluation metrics, outperforming other baselines to a great extent, establishing itself as a versatile and potent framework for interacting with the Windows OS.
We present the detailed breakdown of UFO’s performance across different applications in Table 4 (with the breakdowns for GPT-3.5 and GPT-4 available in the Appendix, Section B). The “-” symbol in the safeguard rate column denotes that all requests related to the application are not sensitive. Notably, UFO demonstrates strong performance across all applications, showcasing its versatility and effectiveness in interacting with and operating on diverse software. However, there is an exception in the case of Adobe Acrobat, where UFO achieves a 60% success rate and a 78.7% completion rate. This can be attributed to the fact that many control types in Adobe Acrobat are not supported by Windows UI Automation, posing challenges for UFO in operating within this application. Importantly, when tasked with completing requests spanning multiple applications, UFO maintains a high level of performance. Despite requiring more steps (average of 9.8) to fulfill such requests, UFO achieves an 80% success rate, an 83% completion rate, and a 100% safeguard rate. This underscores UFO’s sophistication in navigating across different applications to accomplish long-term and complex tasks, solidifying its position as an omnipotent agent for Windows interactions.
4.3 Case Study
To demonstrate the exceptional capabilities of UFO, we provide two case studies illustrating how UFO efficiently fulfills user requests, particularly focusing on tasks involving PowerPoint and spanning across multiple applications [^2] For additional cases, please consult the Appendix Section D..
4.3.1 Deleting All Notes on a PowerPoint Presentation
In Figure 6, we tasked UFO with the request: “Help me quickly remove all notes in the slide of the ufo_testing.”. This request is a common scenario when users want a clean version of a slide without any accompanying notes. Traditional methods involve manually deleting notes page by page, a tedious and time-consuming process for lengthy slides.
However, UFO efficiently identifies a shortcut, simplifying the entire procedure. Analyzing its initial plan, UFO proposes the use of the ‘Remove All Presentation Notes’ function, which is a feature often overlooked by PowerPoint users due to its hidden placement. The automation process commences with UFO navigating to the “File” tab, providing access to the backstage view. Subsequently, it smoothly transitions to the “Info” menu, where the desired function might be located. To examine the document for notes, it clicks the ‘Check for Issues’ button and selects ‘Inspect Document’. Once inspected, the hidden gem, ‘Remove All Presentation Notes’ is revealed at the bottom of the menu. UFO identifies this, scrolls down, locates the function, and initiates the click. Given the sensitive nature of deleting notes, UFO implements its safeguard feature, seeking user confirmation. Upon confirmation, all notes vanish with a single click. This demonstration highlights how UFO empowers users to work smarter, not harder, establishing itself as an intelligent assistant. The video demonstrating this case can be found at the following link: https://github.com/microsoft/UFO/assets/11352048/cf60c643-04f7-4180-9a55-5fb240627834.
4.3.2 Composing an Emails with Information Gathered from Multiple Applications
We make a more complex request shown in Figure 7, “My name is Zac. Please read the meeting note to identify all action items, and include the detailed description LLM training workflow in the LLM-training.png, to compose a new email of these contents. Send the full email to our leader Hidan via email address to ask for his review”. This request puts a significant demand on UFO’s ability to navigate and interact across multiple applications swiftly.
In response, UFO formulates a dynamic plan for the task, deftly navigating between Word, Photos, and Outlook to achieve the overarching objective. Initially, it activates the required document file in Word, utilizing the GetText API to extract text from the main window, which is then logged into its memory. Subsequently, UFO switches to the LLM-training image file in Photos, observing and generating a detailed description for future email composition. With all the necessary information gathered, UFO opens the Outlook application, accessing the “New Email” button to initiate the editing block. The agent then repeats the screenshot capturing, annotation, and action process, autonomously inputting the email recipient, drafting a subject, and composing the email body, including all required information. Prior to sending, the safeguard feature prompts the user for confirmation due to the sensitive nature of the action. Once confirmed, the email is sent. We show the email composed by UFO in Appendix Section C.
The result is a meticulously composed and dispatched email by UFO, skillfully capturing information from both the meeting notes and the details of the LLM pipeline image. This illustrates the remarkable ability of UFO in completing complex tasks that require long-term planning and memory, even when spanning multiple applications. The video demonstrating this case can be found at the following link: https://github.com/microsoft/UFO/assets/11352048/aa41ad47-fae7-4334-8e0b-ba71c4fc32e0.
5 Limitations & Lessons Learned
We acknowledge several limitations in the current UFO framework. Firstly, the available UI controls and actions are currently limited by those supported by pywinauto and Windows UI Automation. Applications and controls that deviate from this standard and backend are not currently supported by UFO. To broaden UFO’s capabilities, we plan to expand its scope by supporting alternative backends, such as Win32 API, or incorporating dedicated GUI models for visual detection, as demonstrated by CogAgent Hong et al. (2023b). This enhancement will enable UFO to operate across a broader range of applications and handle more complex actions.
Secondly, we recognize the challenge UFO faces when exploring unfamiliar application UIs, which may be niche or uncommon. In such cases, UFO may require substantial time to navigate and identify the correct action. To address this, we propose leveraging knowledge from online search engines as an external knowledge base for UFO. Analyzing both textual and image-based guidelines in search results will empower UFO to distill a more precise and detailed plan for completing requests on unfamiliar applications, enhancing its adaptability and generality.
6 Conclusion
We introduce UFO [Uncaptioned image], an innovative UI-focused agent designed to fulfill user requests in natural language through intelligent interactions with applications on the Windows OS. Leveraging GPT-Vision, UFO analyzes application GUI screenshots and controls information to dynamically select the optimal application and control for executing actions and satisfying user queries. The grounding of actions on the application is facilitated by a control interaction module, minimizing human intervention and enabling full automation, establishing it as a comprehensive LAM framework. With the dual-agent design, namely, HostAgent and AppAgent, UFO intelligently switches between applications, allowing for the completion of long and complex tasks that span different applications. Additionally, we incorporate key features such as Action Customization and Safeguard to enhance UFO’s extendibility and safety. Evaluation results based on 50 requests across 9 popular Windows OS applications demonstrate its remarkable versatility and generalizability. To the best of our knowledge, UFO represents the pioneering effort in developing a UI automation agent specifically tailored for the Windows OS environment.
Figure 6: An detailed example of UFO completing user request: “Help me quickly remove all notes in the slide of the ufo_testing.”.
Figure 7: A detailed example of UFO completing user request: “My name is Zac. Please read the meeting note to identify all action items, and include the detailed description LLM training workflow in the LLM-training.png, to compose an new email of these contents. Sent the full email to our leader Hidan via {email address} to ask for his review.”.
Appendix A Requests in WindoesBench and Detailed Evaluations
In Table 5, 6, and 7, we present the complete user requests included in WindowsBench, along with the detailed results achieved by UFO. These requests span across nine different popular Windows applications, incorporating various commonly used functions. Requests with follow-up tasks are sequentially numbered for clarity. In the Safeguard column, “-” denotes that the request is not sensitive and does not require user confirmation. “” indicates that the request is sensitive, and UFO successfully activates the safeguard for user confirmation, while “” signifies that the safeguard fails to trigger for a sensitive request. In the success column, “” denotes that UFO completes the request successfully, while “” indicates a failure to fulfill the request.
Table 5: The requests in WindowsBench and detailed results achieved by UFO (Part-I).
Request Application Safeguard Step Success Completion Rate
I am Zac. Draft an email, to [email address 1] and cc [email address 2], to thanks for his contribution on the VLDB paper on using diffusion models for time series anomaly detection. Don’t send it out. Outlook - 6 100%
Search ’Spring Festival’ in my mail box, and open the second returned email. Outlook - 4 100%
Delete the first junk email. Outlook 4 100%
Please download and save the pdf attachment to the local in the first sent email. Outlook - 10 70%
(1) Draft an email, to {email address 1} and to ask him if he can come to the meeting at 3:00 pm. Don’t send it out. (2) Add some content to the email body to tell him the meeting is very important, and cc {email address 2} as well. (3) Send the email now. Outlook 10 100%
Give a detailed description of the components and workflows of the TaskWeaver architecture in the image. Photos - 2 100%
Open the image of the LLM training and rotate it for 2 times. Photos - 3 100%
Open the image of the LLM training and delete it. Photos 4 100%
Open the mouse image, then loop over the next two images, without closing them. Summarize them one by one. Photos - 5 83.3%
(1) Zoom in the image of the autogen for 2 times to make it clearer, then describe it with detailed workflow and components. (2) It looks too big. Just zoom it to fit the screen. Photos - 6 100%
Help me quickly remove all notes in the slide of the ufo_testing, without looping through each slide one-by-one. PowerPoint 8 100%
Clear the recording on current page for ppt. PowerPoint 3 100%
Please add Morph transition to the current page and the next page of the ufo_testing.ppt PowerPoint - 6 83.3%
Summarize scripts for all pages in the ppt one by one to help with my presetation. PowerPoint - 7 85.7%
(1) Apply the first format generated by the Designer to the current ppt. (2) This format is not beautiful. Select a better format and explain why you select it. PowerPoint - 8 75%
Check spelling and grammar of the current meeting note. Word - 2 100%
Please change the theme of the meeting note into ’Organic’. Word - 4 100%
Save meeting note as an Adobe PDF to the local. Word - 6 83.3%
Add a cover page to the meeting note. Choose one you think is beautiful. Word - 5 100%
(1) Please change to a better-looking format and color to beautify the current page of the meeting note. (2) Do more to make it look even better. Word - 10 80%
Table 6: The requests in WindowsBench and detailed results achieved by UFO (Part-II).
Request Application Safeguard Step Success Completion Rate
Close all pdf that are currently open. Adobe Acrobat 3 100%
Cascade the current pdf windows and close the first one. Adobe Acrobat 3 60%
Visually find and present to me the timeline and milestones in the current meeting note. Adobe Acrobat - 4 50%
(1) Read the imdiffuion paper. What is the main contribution of it? (2) Find the first figure of the paper. What does it show? Adobe Acrobat - 5 100%
(1) Comprehend the current pdf, what is the difference between phase 1 and phase 2? (2) What is the difference of the Application Selection Agent and the Action Selection Agent? (3) Base on the workflow pdf, is there anything can be improved for the agents? (4) Thanks, this is actually the workflow of yourself. Do you think there are other things can be improved for you? I will implement your thought on it. (5) Great. Do you think there are other things that can be improved visually for the figure itself in the pdf, to provide better illustration for other readers? Adobe Acrobat - 6 83.3%
Navigate to /Desktop/mix/screen and open s0.png by only double clicking. File Explorer - 6 83.3%
Delete all files individually in the current figure folder. File Explorer 6 100%
Copy the actagent file in the folder to the Document folder. File Explorer - 5 60%
Create a new txt file in the current folder. File Explorer - 3 100%
(1) Open all images that are related to dragon in the test folder. (2) Summarize in detailed what you have seen in the image d4. File Explorer - 4 100%
Tell me the changelog of the {code base name} repo in their readme base on your observation. Visual Studio Code - 5 80%
Find the use of ’print_with_color’ in the script folder in the {code base name} repo. Visual Studio Code - 4 100%
Download the Docker extension in the {code base name} repo. Visual Studio Code - 4 100%
Create a new terminal at the {code base name} repo. Visual Studio Code - 2 100%
(1) Read through carefully and review the code in the utils.py in the {code base name} and identify any potential bugs or aspects can be improved. (2) What about in the draw_bbox_multi function? Visual Studio Code - 5 40%
Please send a ’smile’ emoticon to the current chatbox at Wechat. WeChat 4 100%
Delete the current chatbox of ’File Transfer’ on Wechat. WeChat 3 100%
Please like {user name}’s post at Moment. WeChat - 3 100%
Bring the first picture to the front in my ”favourite” at WeChat and decribe it. WeChat - 5 100%
(1) Open the {account name} Official Account at WeChat. (2) Send a message to ’File Transfer’ to introduce this account. WeChat 10 90%
Table 7: The requests in WindowsBench and detailed results achieved by UFO (Part-III).
Request Application Safeguard Step Success Completion Rate
How many total citations does Geoffrey Hinton have currently? Edge Browser - 4 100%
Post ”It’s a good day.” on my Twitter. Edge Browser 6 100%
Download the Imdiffusion repo as zip. Edge Browser - 7 100%
Change the theme of the browser to the icy mint. Edge Browser - 5 100%
(1) Find and navigate to {name}’s homepage at Microsoft. (2) Print this page in color. Edge Browser - 5 60%
My name is Zac. Please read the meeting note in meeting_note.docx to identify all action items, and include the detailed description LLM training workflow in the LLM-training.png, to compose an new email of these contents. Sent the full email to our leader Hidan via {email address} to ask for his review. Word, Photos, Outlook - 10 100%
Search for and read through the latest news about Microsoft and summarize it, then send it to ’File Transfer’ on WeChat. Edge Browser, WeChat 9 100%
Open the image of ufo_rv at UFO-windows/logo/, summarize its content and use the summary to search for a similar image on Google. File Explorer, Edge Browser - 10 90%
Observe the overview.pdf, then download the most similar paper base on this architecture from the Internet base on your understanding, and sent it to the ’File Transfer’ on WeChat. Adobe Acrobat, Edge Browser, WeChat - 10 25%
(1) Read the paper.pptx, summarize the paper title presented on the slide, then search and open the paper on the Internet. (2) Can you summarize this paper and download its PDF version for me? PowerPoint, Edge Broswer - 10 100%
Appendix B Performance Breakdown of GPT-3.5 and GPT-4
In Tables 8 and 9, we present the detailed performance breakdown for GPT-3.5 and GPT-4, respectively. It is evident that GPT-4 significantly outperforms GPT-3.5, aligning with our expectations. Notably, GPT-4 exhibits superior performance, particularly in tasks related to Visual Studio Code, where GPT-3.5 frequently struggles to select the correct application initially, leading to its overall lower performance. Another noteworthy instance is seen in Edge Browser, where GPT-3.5 often overlooks crucial steps, resulting in request failures. In general, both baselines exhibit inconsistent performance across different applications, and their overall efficacy falls significantly short when compared to UFO.
Table 8: The detailed performance breakdown achieved by GPT-3.5 (Human Surrogate) on WindowsBench.
Application Success Step Completion Rate Safeguard Rate
Outlook 60.0% 7.6 76.2% 100.0%
Photos 60.0% 5.6 35.7% 100.0%
PowerPoint 20.0% 11.2 20.6% 0.0%
Word 20.0% 7.6 25.0% -
Adobe Acrobat 0.0% 7.4 5.3% 50.0%
File Explorer 40.0% 7.0 53.3% 0.0%
Visual Studio Code 0.0% 7.6 0.0% -
WeChat 40.0% 6.0 48.4% 33.3%
Edge Browser 0.0% 7.6 32.0% 100.0%
Cross-Application 0.0% 11.0 18.8% 50.0%
Table 9: The detailed performance breakdown achieved by GPT-4 (Human Surrogate) on WindowsBench.
Application Success Step Completion Rate Safeguard Rate
Outlook 100.0% 8.4 73.9% 0.0%
Photos 40.0% 7.0 32.7% 100.0%
PowerPoint 40.0% 10.4 35.2% 50.0%
Word 20.0% 9.2 15.3% -
Adobe Acrobat 0.0% 7.6 40.2% 50.0%
File Explorer 80.0% 6.2 63.4% 100.0%
Visual Studio Code 40.0% 7.4 40.3% -
WeChat 40.0% 6.2 68.0% 66.7%
Edge Browser 60.0% 8.2 58.8% 100.0%
Cross-Application 0.0% 13.8 49.7% 50.0%
Appendix C Email Composed by UFO in Section 4.3.2
Below is the email composed by UFO, as detailed in Section 4.3.2. The content of the email is synthesized by gathering information from a Word document and incorporating a description extracted from an image for the writing.
Title: Review of LLM Training Project Action Items and Workflow
Dear Hidan, Please find below the action items identified from our recent meeting, as well as a detailed description of the LLM training workflow. Action Items: - Zac Yang: Develop a comprehensive project plan and share it with the team by 2024.03.06. - Kevin Cheung: Conduct thorough research on potential risks associated with LLM training and present findings by 2024.04.01. - Tom James: Formulate a detailed proposal for resource allocation, including computing resources and software tools. Submit it for review by 2024.01.03. - Sophia Chow: Compile and distribute meeting minutes, including action items and deadlines, within 24 hours. LLM Training Workflow Description: The LLM training framework consists of four main components: ’Pre-training’ with no human involvement using proprietary and unlabeled data, ’Self-supervised learning’ where the base model is trained with narrow datasets, ’Fine-tuning’ which involves human feedback to fine-tune the model, and ’Prompt-based learning’ where the augmented model is developed through prompting with human involvement and specialized knowledge. Please review the above information and provide your feedback at your earliest convenience. Best regards, Zac
Appendix D Additional Case Study
In this section, we present six additional case studies to illustrate the efficacy of UFO in performing diverse tasks across the Windows operating system. These case studies encompass both single-application scenarios and instances where UFO seamlessly transitions between multiple applications.
D.1 Reading a PDF
Figure 8: UFO completes user request: “(1) Read the imdiffuion paper. What is the main contribution of it? (2) Find the first figure of the paper. What does it show? ”
UFO also demonstrates its effectiveness in reading and comprehending a PDF file. In Figure 8, we request UFO to “(1) Read the imdiffusion paper. What is the main contribution of it? (2) Find the first figure of the paper. What does it show?”. The imdiffusion paper Chen et al. (2023c) delves into complex technology related to anomaly detection and diffusion models. Remarkably, UFO showcases its understanding of the paper by summarizing its abstract using OCR in vision input, accurately capturing its main contribution. In a follow-up request, we ask UFO to find and explain the first figure. UFO navigates through the PDF, locates the initial figure, and delivers a detailed explanation, comparing the forecasting, reconstruction, and imputation approaches in anomaly detection. This showcases the remarkable OCR ability of UFO in understanding both text and figures, making it a valuable assistant for paper reading. Below shows UFO’s summary on the paper and the description of the first figure:
Paper Summary: The ’ImDiffusion’ paper’s abstract discusses the importance of anomaly detection in multivariate time series data and the challenges associated with existing approaches. The paper introduces ’ImDiffusion’, a novel approach that combines time series imputation and diffusion models to achieve accurate and robust anomaly detection. The method is designed to handle complex multivariate time series data, enabling the imputation of missing and irregular values in the time series. This approach reportedly improves the accuracy of anomaly detection by leveraging diffusion models as a preprocessing step to accurately reconstruct the anomaly detection time series.
Figure Description: Figure 1: Examples of reconstruction, forecasting and imputation modeling of time series for anomaly detection. To address the challenges and overcome the limitations of existing approaches, we propose a novel anomaly detection method - ImDiffusion. This detector combines the use of time series imputation and diffusion models to achieve accurate and robust anomaly detection. ImDiffusion employs dedicated gating data that makes the diffusion models extra cautiously model the time points, helping utilities see more detail, to recognize valued and serious issues of anomalies through the challenges of detection. The paper introduces a novel framework based on the combination of time series imputation and diffusion models that effectively leverages neighboring values in the time series as additional conditional information, enabling a more accurate modeling of the temporal and inter-related dependencies present in multivariate data. Secondly, the reference performance threshold of the calibration test of the counter prediction from then exchanging values helps to focus the detection in outliers, Figure 1 presents an example in which forecasting, reconstruction, and imputation models are employed to predict a time series using diffusion models. Observe that the imputation-based methods exhibit the highest prediction performance especially on normal periods, thus only the imputation method successfully identifies the period of anomaly. We therefore employ time series imputation for accurate self-supervised modeling of time series, which forms the foundation of our proposed ImDiffusion framework.
D.2 Designing a PowerPoint Slide
Figure 9: UFO completes user request: “(1) Apply the first format generated by the Designer to the current ppt. (2) This format is not beautiful. Select a better format and explain why you select it.”
In Figure 9, we present another example where we make two requests to UFO: “(1) Apply the first format generated by the Designer to the current ppt. (2) This format is not beautiful. Select a better format and explain why you selected it.” The first step is relatively straightforward, where UFO successfully navigates to the ‘Designer’ Pane and applies the desired format to the slide. The second request, however, depends on UFO’s aesthetic judgment. UFO performs this task effectively and provides a reasonable explanation: “I have selected the design suggestion because it provides a clear visual hierarchy by placing the descriptive picture at the top, drawing immediate attention. The title and text are positioned at the bottom, making it easy for the audience to follow the flow of information. The contrasting background colors and font colors enhance readability and focus on the content”. This showcases the remarkable visual ability of UFO, as it can complete open-ended requests based on its own aesthetic sense.
D.3 Downloading an Extension for VSCode
Figure 10: UFO completes user request: “Download the Docker extension in the repo.”.
In Figure 10, we present an example where UFO effectively fulfills a user request in Visual Studio Code: “Download the Docker extension in the repo.”. Although this is a relatively simple task, we aimed to test UFO’s performance in operating on a dark mode GUI. Remarkably, UFO completes this request effortlessly by clicking the Extension button, inputting “Docker” into the search box, and accurately clicking the “Install” button for the Docker Extension. With just one click, the task is successfully completed. This showcases UFO’s capability to operate on less popular applications, even when they are in dark mode.
D.4 Post a Twitter
Figure 11: UFO completes user request: “Post ‘It’s a good day.’ on my Twitter. ”.
In Figure 11, we shift our focus to the Edge Web browser to assess UFO’s capability in operating over one of the most popular application types on Windows OS. Specifically, we request UFO to perform the task of “Post ‘It’s a good day.’ on my Twitter.”, a common yet intricate request for an agent. The observation reveals UFO’s seamless execution, where it inputs the Twitter address into the browser’s address bar, navigates to the link perceived as the main Twitter page, identifies the Post button on the Twitter page, and clicks it. This decision, although not necessarily straightforward, allows UFO to input the required text for the tweet. The safeguard activates before sending, prompting user confirmation. The successful posting of the tweet demonstrates UFO’s sophisticated ability to operate on web browsers, a category of applications widely used on Windows OS.
D.5 Sending the News
Figure 12: UFO completes user request: “Search for and read through the latest news about Microsoft and summarize it, then send it to ’File Transfer’ on WeChat.”.
In a cross-application example, we illustrate how UFO can gather news from the Internet and share it on WeChat in Figure 12. Specifically, we issue the command: “Search for and read through the latest news about Microsoft and summarize it, then send it to ’File Transfer’ on WeChat”. UFO adeptly inputs the query “latest news about Microsoft” into the Google search bar on Edge Browser, initiates the search, and opens the first link in the search results. The page contains multiple news panes, and UFO skillfully summarizes the content using its visual OCR ability, logging the information into its memory. Subsequently, UFO opens WeChat, locates the designated chatbox, and inputs the summarized news retrieved from its memory. Upon user confirmation triggered by the safeguard, UFO sends the news. We show the news sent below.
The Microsoft News page displays several articles. The headlines include: ’Excited to announce the general availability of Copilot for Sales and Copilot for Service, as we continue to extend Copilot to every role and function.’, ’Microsoft announces quarterly earnings release date’, ’With Copilot Pro, we’re helping even more people supercharge their creativity and productivity by unlocking Copilot in Microsoft 365 apps, providing access to the very latest models 2014 and more.’, and ’Microsoft unveils new generative AI and data solutions across the shopper journey, offering copilot experiences through Microsoft Cloud for Retail’.
This example once again underscores UFO’s ability to seamlessly transition between different applications, allowing it to effectively and safely complete long-term, complex tasks. Such demonstrations position UFO as an advanced and compelling agent for Windows OS.
D.6 Search Paper in a PowerPoint Slide and Summarize
Figure 13: UFO completes user request: “(1) Read the paper.pptx, summarize the paper title presented on the slide, then search and open the paper on the Internet. (2) Can you summarize this paper and download its PDF version for me?”.
Lastly, in Figure 13, we present an intricate cross-application example with multiple sub-requests where UFO is tasked with finding a paper presented in a slide on the Internet, summarizing the paper, and downloading it. The user’s request is as follows: “(1) Read the paper.pptx, summarize the paper title presented on the slide, then search and open the paper on the Internet. (2) Can you summarize this paper and download its PDF version for me?” This request involves multiple instances of OCR usage by UFO.
To fulfill the first sub-request, UFO carefully observes the title page of the slide, successfully identifies the paper titled ”Xpert: Empowering Incident Management with Query Recommendations via Large Language Models.” Subsequently, it opens a Google page in the Edge Browser, inputs the title into the search bar, and locates the paper as the second result by skipping the first ad link. UFO clicks and opens the link on arXiv, completing the first request.
Upon receiving the second sub-request to summarize the paper and download the PDF version, UFO clicks the “Download as PDF” option and uses its OCR ability to generate the following summary:
The paper titled ‘Xpert: Empowering Incident Management with Query Recommendations via Large Language Models’ discusses the role of large-scale cloud systems in modern IT infrastructure and the impact of incidents occurring within these systems. The paper introduces Xpert, a machine learning framework that automates Kusto Query Language (KQL) recommendations, tailored to new incidents. By leveraging historical incident data and large language models, Xpert generates customized KQL queries, enabling a thorough evaluation of query quality from three comprehensive perspectives. The paper details the deployment of Xpert in the real production environment of a large-scale incident management system in Microsoft, demonstrating its effectiveness in supporting incident management.
This summary aligns perfectly with the paper’s contribution as outlined in Jiang et al. (2023), showcasing UFO’s excellent ability to operate across different applications and its remarkable OCR capabilities.